# Performance Analysis of NR Sidelink and Wi-Fi Coexistence Networks in Unlicensed Spectrum

### v1

@@ III Collaborative Channel Access (CCHA)
The 3GPP standard mandates that the minimum resource allocation unit for the PSSCH in the time domain is a slot. As a result, SL-U users operating in the unlicensed spectrum cannot immediately begin transmission after completing the LBT process; instead, they must wait until the next SSB to transmit data. This limitation significantly impacts the performance of SL-U systems. In a distributed resource allocation scheme, namely Mode 2, each SL-U user independently performs the LBT process and waits for the next SSB to transmit. However, this approach often results in frequent and severe collisions, as illustrated in Fig. 3LABEL:sub@fig:CCHA_1. On the other hand, if SL-U users adopt a centralized resource allocation scheme, the collision issue can be mitigated. Nevertheless, if only Type 1 LBT is employed, the LBT process may fail due to the channel occupied in the previous slot by other SL-U users, resulting in low channel resource utilization, as shown in Fig. 3LABEL:sub@fig:CCHA_2. To overcome these challenges, we propose a collaborative channel access mechanism that optimizes resource utilization and improves the overall performance of SL-U systems.
@@
The mechanism integrates multiple LBT types with a centralized resource allocation strategy to enhance efficiency and coordination. First, the BS performs the Type 1 LBT process to compete for channel access. Upon successful completion, the BS obtains the channel’s time-domain occupancy period, denoted as $T_{cot}$, and the channel bandwidth $B$ in the frequency domain. The BS then centrally allocates these resources, dividing them into slots for SL-U users with data to transmit within its coverage area. SL-U users assigned specific channel resources validate their availability by performing the Type 2 LBT process at a designated position before the SSB. If the Type 2 LBT process is completed, the user proceeds with data transmission, as illustrated in Fig. 3LABEL:sub@fig:CCHA_3.
@@
For LBT channel access, this study specifies that all channel sensing processes must begin at the SSB for experimental convenience, except for specific channel sensing positions designated by the base station. While implementing the collaborative channel access mechanism proposed in this study, the base station and SL-U users exchange control signals to facilitate collaborative operations. To ensure the reliability of control signaling transmission, these signals are conveyed over the licensed spectrum. Furthermore, by the standard, SL-U users avoid transmitting data during the final orthogonal frequency division multiplexing (OFDM) symbol of the allocated resources to establish a guard interval (GI).
@@ IV Joint Optimization of Channel Access and Power Control based on C-GHDRL
In this section, we first elaborate on the fundamental concepts of semi-Markov decision processes and hierarchical reinforcement learning. Subsequently, we introduce the C-GHDRL algorithm, designed to achieve the joint optimization of channel access and power control.
@@
The joint optimization of channel access and power control can be decomposed into two sub-problems: channel access and power control. This hierarchical structure with multi-step decomposition is particularly well-suited for the GHDRL algorithm. However, in the unlicensed spectrum, because of the absence of an information-sharing mechanism among SL-U users, SL-U users are restricted to sense only their local environment. This restriction prevents SL-U users from obtaining complete global information, which hinders the GHDRL algorithm model from making globally optimal decisions.
@@
In the framework, the C-GHDRL algorithm comprises two components: a meta-controller and a controller. The meta-controller is responsible for optimizing channel access, while the controller focuses on power control, utilizing the state information and outputs provided by the meta-controller. Since the state transition probabilities are unknown, we employ a deep Q-learning algorithm to train the policies of both the meta-controller and the controller within the C-GHDRL frame [41]. Specifically, the controller estimates the state-value function $Q_{1}(s,a;g)$ as follows:
@@
where $g$ represents the goal of agent in state $s$ and $\pi_{ag}=\mathcal{P}(a|s,g)$ is the action policy. $r_{t}$ represents the internal reward provided to the controller.
Similarly, for the meta controller, the state-value function $Q_{2}(s,g)$ is defined as:
@@
1:  Initialize experience replay memories $\{\mathcal{D}_{1},\mathcal{D}_{2}\}$ to capacity $\{\mathcal{C}_{1},\mathcal{C}_{2}\}$ for the controller and meta-controller respectively
2:  Initialize $\{Q_{1},Q_{2}\}$ network parameters with random weights $\{\theta_{1},\theta_{2}\}$ for the controller and meta-controller respectively
3:  Initialize target $\{Q^{\prime}_{1},Q^{\prime}_{2}\}$ network parameters with random weights $\{\theta_{1}^{{}^{\prime}},\theta_{2}^{{}^{\prime}}\}$ for the controller and meta-controller respectively
4:  Initialize exploration probability $\varepsilon_{1,g}=1$ for the controller for all goals $g$ and $\varepsilon_{2}=1$ for the meta-controller.
@@
13:           Execute $a$ and obtain next state $s^{\prime}$ and obtain intrinsic reward $r(s,a,s^{\prime})$ from controller
@@
2) Subgoal: The subgoal space $\mathcal{G}$ output by the upper layer meta controller is defined as follows: if the channel is to be accessed, $g(t)=1$; otherwise, $g(t)=0$.
3) State: For the upper layer meta controller, its state is
@@
where $p_{s}^{T}(t-1)$ represents the power sent by the SL-U user at the last slot, and $\boldsymbol{p}_{ch}^{R}(t)$ represents the channel energy detected by all SL-U users except itself. For the lower layer controller, it is an extended state represented as follows:
@@
We use a deep Q network (DQN) to learn optimal policies for the controller and meta controller. Algorithm 1 shows the specific steps of our C-GHDQN method.
@@ V Simulation Result
We simulate a coexistence scenario in which SL-U and Wi-Fi users operate within a $400\times 400\ m^{2}$ indoor environment. In this simulation, a transmission from either SL-U or Wi-Fi users is considered successful when the conditions $\eta_{s,m}\geq\eta_{s,min}$ and $\eta_{w,n}\geq\eta_{w,min}$ are met. The traffic model for users in the coexistence system follows FTP Model 3 [41]. For clarity and convenience, the detailed simulation parameters are summarized in Table II.
@@
To evaluate the effectiveness of our proposed CCHA mechanism, we analyze the experimental results from four comparative schemes. In these simulations, we set the transmission power for both SL-U users and Wi-Fi users at 23 dBm. Furthermore, we established a lower channel sensing threshold to reduce the possibility of resource reuse between SL-U users and Wi-Fi users, thereby highlighting the advantages of the CCHA mechanism.
@@
We compare the differences between the scheme proposed in this paper and other schemes in terms of Jain’s fairness index, packet reception rate (PRR), throughput, and waiting time. PRR is the ratio of the number of successfully received packets to the total number of packets sent. The waiting time is the duration from the generation to the successful transmission of a transport block. Throughput, denoted by $Th$, is defined as the rate of network transmission per unit of time, with the unit expressed in bits per second (bps).
@@
Fig. 5 illustrates the comparison of Jain’s fairness index across the four schemes as the number of system users increases. First, it can be seen from the figure that as the number of users increases, the Jain’s fairness index of all schemes decreases. The decrease is caused by system resources becoming increasingly scarce as the number of users increases. Then, we observed that the scheme utilizing the CCHA is superior to other comparative schemes regarding Jain’s fairness index. The reasons include two aspects. On the one hand, the advantages of the two types of channel access are fully utilized in the CCHA mechanism. Type 1 LBT performed by the base station can sense whether the channel is idle initially, while Type 2 LBT performed by the SL-U user can sense whether the channel is idle again before SSB, with a shorter sensing time compared to Type 1 LBT. Therefore, the mechanism effectively ensures fairness while enhancing resource utilization. On the other hand, centralized resource allocation (CRA) is used in the CCH, which can ensure that the resources in the SL-U system are completely orthogonal. This method effectively prevents collisions arising from multiple SL-U users selecting the same resources for data transmission due to the absence of a coordination mechanism in the DRA. In addition, the scheme of T1O-LBT and DRA exhibits the poorest performance in terms of Jain’s fairness index. This is because the T1O-LBT scheme requires SL-U users to perform only Type 1 LBT without reconfirming whether the channel is idle before access. This limitation significantly increases the likelihood of collisions with users already occupying the channel. Furthermore, the absence of a coordination mechanism in the DRA scheme leads to increased potential for collisions among SL-U users, further degrading system performance.
Figure 6: Comparisons of PRR over difference schemes.
Figure 7: Comparisons throught over difference schemes.
Fig. 6 and Fig. 7 illustrate the PRR and throughput results for varying numbers of users in the system. The results indicate that PRR and throughput degrade as the number of users increases. This performance decline is attributed to the increasing scarcity of resources as the user density rises. Furthermore, if the number of users increases to a certain scale, the performance deteriorates sharply due to severe resource contention. Notably, under moderate user density, the CCHA scheme outperforms others regarding PRR and throughput. This improvement is due to the CCHA scheme’s ability to effectively mitigate the conflict inherent in distributed resource allocation methods while enhancing channel access opportunities for SL-U users in the unlicensed spectrum.
Fig. 8 presents the waiting time for SL-U and Wi-Fi users under different channel access mechanisms and varying numbers of system users. Since the channel access mechanism is the primary factor influencing waiting time, we focus on comparing the waiting time across different channel access types with centralized resource allocation. First, as the number of users increases, we can observe that the waiting time also increases. The reason is that increased competition for access channels intensifies channel contention, leading to longer delays. The waiting time under the CCHA scheme is shorter than that under the CCHA-T1 scheme, as in Fig. 8LABEL:sub@fig:sub.1. This difference arises because, in the CCHA mechanism, SL-U users perform Type 2 LBT before the SSB. This allows the channel status to be determined in a shorter sensing period, enabling SL-U users to access the channel more quickly than those performing Type 1 LBT under the CCHA-T1 scheme. Conversely, as shown in Fig. 8LABEL:sub@fig:sub.2, the waiting time for Wi-Fi users in a system employing the CCHA mechanism is slightly longer than that for Wi-Fi users in a system using the CCHA-T1 mechanism, which indicates that, while the CCHA mechanism enhances access opportunities for SL-U users, it introduces a certain degree of impact on Wi-Fi users. However, this impact remains acceptable, ensuring effective coexistence between SL-U and Wi-Fi users. Furthermore, as the number of users increases beyond a certain scale, the waiting times for users under both schemes are relatively close, indicating that the CCHA scheme is not suitable for scenarios involving large-scale user deployments.
(a) SL-U Users
(b) Wi-Fi Users
Figure 8: Comparisons waiting time between CCHA and CCHA-T1.
@@
As the number of users increases, system performance degrades significantly due to resource constraints. To address this, SL-U and Wi-Fi users are allowed to perform appropriate resource reuse, improving the overall performance of the coexistence system. To achieve this, we employ the C-GHDRL algorithm to jointly optimize the channel access and transmission power for SL-U users, further enhancing system performance. In this section, the numbers of SL-U and Wi-Fi users are set to 32 and 44, respectively. Additionally, the transmission power of SL-U users is set between $-40$ dBm and 0 dBm due to the relatively short distance between their transmitters and receivers. Meanwhile, Wi-Fi users dynamically adjust their transmission power using a closed-loop power control strategy to adapt to the dynamic variations in the channel environment.
To verify the performance of the C-GHDRL algorithm, we utilize four schemes for comparison, i.e., the DQN algorithm, random algorithm, open-loop power control algorithm, and the fundamental scheme of CCHA (F-CCHA).
@@
    DQN algorithm: The DQN algorithm-based scheme builds upon the F-CCHA scheme and leverages the DQN algorithm to dynamically adjust transmission power for SL-U users.
@@
    Random algorithm: The random algorithm-based scheme builds upon the F-CCHA scheme and leverages the random algorithm to dynamically adjust channel access and transmission power for SL-U users.
@@
    Open-loop power control algorithm: The open-loop power control algorithm-based scheme builds upon the F-CCHA scheme and leverages the open-loop power control algorithm to dynamically adjust transmission power for SL-U users.
@@
Figure 10: Comparisons the PRR over difference schemes.
Figure 11: Comparisons the throught over difference schemes.

### v2

@@ III Collaborative Channel Access (CCHA)
SL-U users operating in the unlicensed spectrum cannot immediately begin transmission after completing the LBT process; instead, they must wait until the next spectrum slot boundary (SSB) to transmit data. This limitation significantly impacts the performance of SL-U systems. In a distributed resource allocation scheme, namely Mode 2, each SL-U user independently performs the LBT process and waits for the next SSB to transmit. However, this approach often results in frequent and severe collisions, as illustrated in Fig. 3LABEL:sub@fig:CCHA_1. In the figure, SL-U users A and B independently perform Type 1 LBT channel sensing at specified locations. Although they sense the channel for different durations, they both wait for the next SSB to transmit data. And, since users A and B cannot know each other’s channel occupancy status in the distributed system, they transmit data in the same slot, resulting in collision.
On the other hand, if SL-U users adopt a centralized resource allocation scheme, the collision issue can be mitigated. Nevertheless, if only Type 1 LBT is employed, the LBT process may fail due to the channel occupied in the previous slot by other SL-U users, resulting in low channel resource utilization, as shown in Fig. 3LABEL:sub@fig:CCHA_2. In Fig. 3LABEL:sub@fig:CCHA_2, SL-U users A, B, C, and D are assigned defined slots to transmit data. To comply with unlicensed spectrum occupancy rules, users must sense the channel at specified locations. Since user A occupies the channel, user B senses the channel as busy during channel sensing, preventing user B from transmitting data in the defined slot. The same applies to user D. This results in low resource utilization.
@@
To address these issues, we propose a collaborative channel access mechanism designed to optimize resource utilization and enhance the overall performance of coexistence systems. This mechanism integrates multiple LBT types with a centralized resource allocation strategy to improve efficiency and coordination. First, the BS performs the Type 1 LBT process to compete for channel access. Upon successful completion, the BS obtains the channel’s time-domain occupation period, denoted as $T_{cot}$, and the channel bandwidth $B$ in the frequency domain. The BS then centrally allocates these resources, dividing them into slots for SL-U users with data to transmit within its coverage area. SL-U users allocated specific channel resources confirm their availability by performing a Type 2 LBT process at designated locations before the SSB. According to the standard, SL-U users avoid transmitting data during the final orthogonal frequency division multiplexing (OFDM) symbol of the allocated resources to establish a guard interval (GI). For a subcarrier spacing of 30 kHz, this guard interval is 33.33 $\mu s$, which is sufficient to complete the Type 2 LBT process, with a sensing time of at most 25 $\mu s$. If the Type 2 LBT process is completed, the user proceeds with data transmission, as illustrated in Fig. 3LABEL:sub@fig:CCHA_3. The mechanism avoids collisions, as shown in Fig. 3LABEL:sub@fig:CCHA_1, by enabling the BS to allocate resources centrally. Furthermore, by using only the Type 2 LBT process, the channel sensing time is significantly shorter than the Type 1 LBT process, helping to prevent resource wastage, as seen in Fig. 3LABEL:sub@fig:CCHA_2. By improving resource utilization, the mechanism enhances the performance of the coexistence system.
@@
When implementing the collaborative channel access mechanism proposed in this study, the base station and SL-U users exchange control signals to facilitate collaborative operations. To ensure the reliability of control signaling transmission, these signals are conveyed over the licensed spectrum.
@@ IV Joint Optimization of Channel Access and Power Control based on C-GHDRL
In this section, we first present a brief background of semi-Markov decision processes and hierarchical reinforcement learning. Subsequently, we introduce the C-GHDRL algorithm, designed to achieve the joint optimization of channel access and power control.
@@
The joint optimization of channel access and power control can be decomposed into two sub-problems: channel access and power control. This hierarchical structure with multi-step decomposition is particularly well-suited for the GHDRL algorithm. However, in the unlicensed spectrum, SL-U users cannot transmit simultaneously due to competition for channel access, and they are unable to receive timely feedback on the environment, which hinders the GHDRL algorithm model from making globally optimal decisions.
@@
In the framework, the C-GHDRL algorithm comprises two components: a channel access controller and a power control controller. The channel access controller is responsible for optimizing channel access, while the power control controller focuses on power control, utilizing the state information and outputs provided by the channel access controller. Since the state transition probabilities are unknown, we employ a deep Q-learning algorithm to train the policies of both the channel access controller and the power control controller within the C-GHDRL frame [38]. Specifically, the power control controller estimates the state-value function $Q_{1}(s,a;g)$ as follows:
@@
where $g$ represents the goal of agent in state $s$ and $\pi_{ag}=\mathcal{P}(a|s,g)$ is the action policy. $r_{t}$ represents the internal reward provided to the power control controller.
@@
1:  Initialize experience replay memories $\{\mathcal{D}_{1},\mathcal{D}_{2}\}$ to capacity $\{\mathcal{C}_{1},\mathcal{C}_{2}\}$ for the power control controller and channel access controller respectively
2:  Initialize $\{Q_{1},Q_{2}\}$ network parameters with random weights $\{\theta_{1},\theta_{2}\}$ for the power control controller and channel access controller respectively
3:  Initialize target $\{Q^{\prime}_{1},Q^{\prime}_{2}\}$ network parameters with random weights $\{\theta_{1}^{{}^{\prime}},\theta_{2}^{{}^{\prime}}\}$ for the power control controller and channel access controller respectively
4:  Initialize exploration probability $\varepsilon_{1,g}=1$ for the power control controller for all goals $g$ and $\varepsilon_{2}=1$ for the channel access controller.
@@
13:           Execute $a$ and obtain next state $s^{\prime}$ and obtain intrinsic reward $r(s,a,s^{\prime})$ from power control controller
@@
Similarly, for the channel access controller, the state-value function $Q_{2}(s,g)$ is defined as:
@@
2) Subgoal: The subgoal space $\mathcal{G}$ output by the upper layer channel access controller is defined as follows: if the channel is to be accessed, $g(t)=1$; otherwise, $g(t)=0$.
3) State: For the upper layer channel access controller, its state is
@@
where $p_{s}^{T}(t-1)$ represents the power sent by the SL-U user at the last slot, and $\boldsymbol{p}_{ch}^{R}(t)$ represents the channel energy detected by all SL-U users except itself. For the lower layer power control controller, it is an extended state represented as follows:
@@
We use a deep Q network (DQN) to learn optimal policies for the power control controller and channel access controller. Algorithm 1 shows the specific steps of our C-GHDQN method.
@@ V Simulation Result
We simulate a coexistence scenario in which SL-U and Wi-Fi users operate within a $400\times 400\ m^{2}$ environment. In this simulation, a transmission from either SL-U or Wi-Fi users is considered successful when the conditions $\eta_{s,m}\geq\eta_{s,min}$ and $\eta_{w,n}\geq\eta_{w,min}$ are met. The traffic model for users in the coexistence system follows FTP Model 3 [38]. For clarity and convenience, the detailed simulation parameters are summarized in Table II.
@@
To evaluate the effectiveness of our proposed CCHA mechanism, we compare it with three benchmark schemes. Given that the CCHA mechanism involves two key variables, allocation scheme and channel access. To ensure a comprehensive comparative analysis, we employ the control variable method to design the three benchmark schemes.
@@
Figure 5: Comparisons of PRR over difference schemes.
Figure 6: Comparisons throught over difference schemes.
In these simulations, we set the transmission power for both SL-U users and Wi-Fi users at 23 dBm. Furthermore, we established a lower channel sensing threshold to reduce the possibility of resource reuse between SL-U users and Wi-Fi users, thereby highlighting the advantages of the CCHA mechanism.
We compare the differences between the scheme proposed in this paper and other schemes in terms of packet reception rate (PRR), throughput and Jain’s fairness index. PRR is the ratio of the number of successfully received packets to the total number of packets sent. Throughput, denoted by $Th$, is defined as the rate of network transmission per unit of time, with the unit expressed in bits per second (bps).
@@
Fig. 5 illustrates the PRR results for different numbers of users in the system. As shown in the figure, the PRR for all schemes decreases as the number of users increases. The decrease is caused by system resources becoming increasingly scarce as the number of users increases. Furthermore, if the number of users increases to a certain scale, the performance deteriorates sharply due to severe resource contention. Under moderate user density, the CCHA scheme outperforms the other schemes in terms of PRR. The reasons include two aspects. On the one hand, the advantages of the two types of channel access are fully utilized in the CCHA mechanism. Type 1 LBT performed by the base station can sense whether the channel is idle initially, while Type 2 LBT performed by the SL-U user can sense whether the channel is idle again before SSB, with a shorter sensing time compared to Type 1 LBT. Therefore, the mechanism effectively improves the successful acceptance rate of data packets. On the other hand, CCHA uses centralized resource allocation (CRA), which can ensure that the resources in the SL-U system are completely orthogonal. This method effectively prevents collisions arising from multiple SL-U users selecting the same resources for data transmission due to the absence of a coordination mechanism in the DRA, further increasing the likelihood of conflicts and reducing the successful packet acceptance rate. Additionally, the PRR performance of the T1O-LBT and DRA schemes is the worst. The T1O-LBT scheme requires SL-U users to perform only Type 1 LBT without reconfirming whether the channel is idle before access. This limitation greatly increases the possibility of collisions with users who have already occupied the channel, leading to a higher rate of packet transmission failure. Furthermore, the absence of a coordination mechanism in the DRA scheme leads to increased potential for collisions among SL-U users, further diminishing the number of successfully transmitted packets.
Fig. 6 illustrate the throughput results for varying numbers of users in the system. The throughput performance trend in the figure mirrors is similar to the PRR. This is because the CCHA solution enhances resource utilization and mitigates collision issues, resulting in better throughput performance in the coexistence system compared to the three benchmark solutions.
Fig. 7 illustrates the comparison of Jain’s fairness index across the four schemes as the number of system users increases. As shown in the figure, Jain’s fairness index decreases for all schemes as the number of users rises. The decrease is caused by system resources becoming increasingly scarce as the number of users rises. The CCHA scheme outperforms the other comparison schemes regarding the Jain fairness index. This is because the CCHA mechanism leverages the advantages of two types of channel access and centralized resource allocation, effectively ensuring fairness while enhancing resource utilization. In addition, the T1O-LBT and DRA schemes exhibit the worst Jain fairness index performance. This is because SL-U users only perform Type 1 LBT and do not reconfirm the channel status before accessing the channel. This limitation increases the likelihood of collisions with Wi-Fi users, leading to a lower fairness index.
@@
As the number of users increases, system performance degrades significantly due to resource constraints. To address this, SL-U and Wi-Fi users are allowed to perform appropriate resource reuse, improving the overall performance of the coexistence system. To achieve this, we employ the C-GHDRL algorithm to jointly optimize the channel access and transmission power for SL-U users, further enhancing system performance. In this section, the numbers of SL-U and Wi-Fi users are set to 32 and 44, respectively. Additionally, the transmission power of SL-U users is set between $-40$ dBm and 0 dBm due to the relatively short distance between their transmitters and receivers. Meanwhile, Wi-Fi users dynamically adjust their transmission power using a closed-loop power control strategy to adapt to the dynamic variations in the channel environment [39].
To verify the performance of the C-GHDRL algorithm, we utilize four schemes for comparison, i.e., the DQN algorithm, the fundamental scheme of CCHA (F-CCHA), random algorithm, and open-loop power control algorithm.
@@
    DQN algorithm [40]: The DQN algorithm-based scheme builds upon the F-CCHA scheme and leverages the DQN algorithm to dynamically adjust transmission power for SL-U users.
@@
    Random algorithm [41]: The random algorithm-based scheme builds upon the F-CCHA scheme and leverages the random algorithm to dynamically adjust channel access and transmission power for SL-U users.
@@
    Open-loop power control algorithm [42]: The open-loop power control algorithm-based scheme builds upon the F-CCHA scheme and leverages the open-loop power control algorithm to dynamically adjust transmission power for SL-U users.
Figure 8: Comparisons the PRR over difference schemes.
Figure 9: Comparisons the throught over difference schemes.